What is the pOH of a 0.0031M solution of the strong acid HBr?
1
Expert's answer
2020-05-01T14:41:11-0400

HBr "\to" H+ + Cl-

pH = - log[H+] = -log(0.0031)= 2.51

pH + pOH = 14

pOH=14 - pH = 14-2.51= 11.49